Two sides of a triangle are given by the roots of the equation x^2-5x+6=0 and angle between the sides is 60°. Then the perimeter of the triangle is

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
On Solving quadratic equations, we get x = 3 and 2
 
Hence sides are 3 and 2
 
Now
 
Cos 60 = 3² + 2² - c² /(2*3*2)
 
1/2 = 13 -c² /12
 
6 = 13 -c²
 
c² = 7
Hence c = √7
 
Now perimeter = 2 +3 + √7 = 5 + √7